itzsummer:

Valerie cut a piece of wood at a 72° angle for her project. What is the degree measure of the supplementary angle on the leftover piece of wood?

AP:

Supplementary angles add up to 180º So, 72º + x = 180º Solve for x

itzsummer:

x=18?

AP:

Does 72 + 18 = 180?

itzsummer:

yes

AP:

No, 72 + 18 = 90 Not 180...

itzsummer:

ohhhh

itzsummer:

i need to use a calculator wait lol

itzsummer:

108

AP:

Correct